Car A of mass 2000kg moving at 73 km/h tries to overtake car B of mass 1250 kg moving at 60 km/h,east. Car A collides with car B. The cars stick together after collision. Find their velocity after collission.

1
Expert's answer
2022-01-26T15:31:47-0500

Conservation of momentum is the key:


"m_1v_1+m_2v_2=(m_1+m_2)u,\\\space\\

u=(m_1v_1+m_2v_2)/(m_1+m_2)=68\text{ km/h}."